OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

helppppppppppppp:(

OpenStudy (anonymous):

start with \[\cos(3x)=\cos(2x+x)\] then use the "addition angle" formula for cosine

OpenStudy (anonymous):

\[\cos(2x+x)=\cos(2x)\cos(x)-\sin(2x)\sin(x)\] then use the "double angle" formula for sine and cosine to finish

OpenStudy (anonymous):

im at cos 3x= cos (2x+x) = \[(\cos ^{2} x-\sin^2 x) \cos x - (2\sin x \cos x)(\sin x)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have done what is asked your answer is in terms of \(\sin(x)\) and \(\cos(x)\)

OpenStudy (anonymous):

you could maybe clean it up with some algebra

OpenStudy (anonymous):

im not able to clean it up like i dont know how to simplify (2sinx cosx) (sinx )

OpenStudy (anonymous):

\[-2\sin(x)\cos(x)\sin(x)=-2\sin^2(x)\cos(x)\]

OpenStudy (anonymous):

and \[(\cos^2(x)-\sin^2(x))\cos(x)=\cos^3(x)-\sin^2(x)\cos(x)\] via the distributive law

OpenStudy (anonymous):

the answer choices are: cos x - 4 cos x sin^2x -sin^3x + 2 sin x cos x -sin^2x + 2 sin x cos x 2 sin^2x cos x - 2 sin x cos x literally nothign matches